# Geometric theory of constrained Schrödinger dynamics with application to time-dependent density-functional theory on a finite lattice

### v1

@@ II.2 Geometric structure of the set of constrained states for time-independent constraints
We have to be careful to work with the real structure of our state space $\mathbb{C}^{d}$, and not the complex structure. This is because it is not the same to be differentiable in the real sense or in the complex sense. By real structure we mean that we see any vector $\psi$ in $\mathbb{C}^{d}$ as a vector in $\mathbb{R}^{2d}$, each component $\psi_{j}$ being split into its real part $\Re(\psi_{j})$ and imaginary part $\Im(\psi_{j})$. The usual real scalar product in $\mathbb{R}^{2d}$ of two vectors $\psi,\psi^{\prime}\in\mathbb{C}^{d}$ is
@@
The set $\mathcal{C}$ in (4) can be decomposed into a regular part $\mathcal{M}$ and a singular part $\mathcal{S}$ as $\mathcal{C}=\mathcal{M}\cup\mathcal{S}$. The regular part $\mathcal{M}$ is the set where the constraints are non-redundant (or qualified in the language of constrained optimization) in the sense that the $\mathcal{O}_{m}\psi$’s are $\mathbb{R}$-linearly independent. It is easily seen that $\mathcal{M}$ is a smooth (infinitely differentiable) submanifold of $\mathbb{R}^{2d}$ of dimension 2dM2d-M. The singular part $\mathcal{S}$ is interpreted as a boundary of $\mathcal{M}$, where (some of) the constraints become redundant. We will restrict ourselves to states evolving only within the regular part $\mathcal{M}$.
@@
$\mathcal{T}_{\psi}:=\big\{h\in\mathbb{C}^{d}\ :\ \Re{\left\langle h,\mathcal{O}_{m}\psi\right\rangle}=0,\ \forall m=1,...,M\big\}.$ (5)
@@
We assume that our initial state belongs to the regular part $\psi_{0}\in\mathcal{M}$. That is, we assume that $S^{\psi_{0}}$ is invertible. In addition, we will only work with trajectories $t\mapsto\psi(t)$ defined over an interval of times $t\in[0,T)$ for which $S^{\psi(t)}$ remains invertible. Note that a matrix very similar to the matrix $S^{\psi}$ already appeared in a work on the abstract extension of the Hohenberg–Kohn theorem [18]. For the special case of commuting observables $\mathcal{O}_{m}$’s, the matrix $S^{\psi}$ also appears in a recent work [20] in the context of constrained search in imaginary time for ground-state DFT.
@@ II.3 Time-dependent constraints
From the definition (6) of $\mathcal{N}_{\psi}$, we can write
@@
Historically, this problem was studied first by Dirac [21] and Frenkel [22, p. 253]. These authors were however considering the manifold of Slater determinants, that has a natural complex structure, so that several of the difficulties we will encounter do not occur in their case. Meyer, Kǔcar and Cederbaum were probably the first to notice in [23] that several definitions of the constrained Schrödinger dynamics that were the same for Dirac and Frenkel can give different results in other situations. The general theory was further developed in [24, 25, 26, 27, 28] but, to the best of our knowledge, it was never applied to the case of constraining expectation values $\left\langle\psi(t),\mathcal{O}_{m}\psi(t)\right\rangle$}. This is the situation of interest for TDDFT. In the next two sections we describe the main two methods used in practice to choose the correction term $F$ so that the solution to the modified Schrödinger equation stays on the considered manifold at all times, and specify the expressions of $F$ in the case of a manifold defined by such constrained expectation values. We will also introduce a new family of constrained Schrödinger equations that interpolate between the previous two methods, which does not seem to have been considered so far.
@@ III.1 Stationarity of the action
In this section, we investigate under which conditions on the observables $\mathcal{O}_{m}$ one can define a unique constrained dynamics solely based on the stationarity of the action. We emphasize that this is the traditional point of view of TDDFT, as we will recall below, hence it is important to understand how this works (or, in fact, why it does not work so easily in this particular case).
@@
The details of this argument can for instance be read in [28, Prop. 3.1]. Stationarity of the action can therefore be re-interpreted as the requirement that the residue of the unconstrained Schrödinger equation $i\partial_{t}\psi(t)-H(t)\psi(t)$ belongs to the normal space of $\psi(t)$ all along the trajectory. This is to make it orthogonal to the small deformations of the path that can only happen in the directions of the tangent space. We emphasize once again the importance of the real structure of the problem in this result due to the appearance of the real scalar product.
@@ III.3 The case of commuting observables
Exactly as before we note that the invertibility of $K^{\psi(t)}(t)$ implies the full-rank condition (7), hence the stationarity of the action for the trajectory $\psi(t)$ provided by the theorem. Also, the normalization $\|\psi(t)\|=1$ is automatic and the identity matrix should not be included in the list of the $\mathcal{O}_{m}$’s, otherwise the matrix $K^{\psi(t)}(t)$ cannot be invertible.
@@ IV.1 Geometric Schrödinger equation
A manifold for which the tangent and normal spaces are complex-linear and not just real-linear (i.e., $i\mathcal{T}_{\psi}=\mathcal{T}_{\psi}$ and $i\mathcal{N}_{\psi}=\mathcal{N}_{\psi}$) is called a Kähler manifold. For such a manifold, the variational principle (17) and the geometric principle (28) give the same answer, possibly up to an irrelevant global phase factor. When the manifold is defined by fixing expectation values of some observables $\mathcal{O}_{m}$, the spaces $\mathcal{T}_{\psi}$ and $\mathcal{N}_{\psi}$ will usually not be complex-linear, however. For instance, if the $\mathcal{O}_{m}$’s are commuting matrices (as they are in TDDFT) then we have $\Im{\left\langle\mathcal{O}_{m}\psi,\mathcal{O}_{n}\psi\right\rangle}=0$ for all $m,n$, as we have seen in the previous section. This implies $\Re{\left\langle i\mathcal{O}_{m}\psi,\mathcal{O}_{n}\psi\right\rangle}=0$, that is $i\mathcal{N}_{\psi}\subset\mathcal{T}_{\psi}$.
Since the condition (28) is only based on an orthogonal projection relying on the geometric structure of the constraint manifold $\mathcal{M}$, we call it the geometric principle (GP). In the literature, the latter has been given different names. It is often called the McLachlan principle [31]. The interpretation in terms of a projection on the tangent space was already mentioned by Frenkel in [22, p. 253] and, apparently, in an appendix of the Russian version of the book [21] by Dirac (see the comments in [32]). The McLachlan principle was rediscovered by Nazarov in [33, 34].
@@
Because of the $i$, at first sight it may seem that we are introducing a non-Hermitian perturbation of the Hamiltonian $H(t)$. But this is not the best way to see it. Recall that this is a nonlinear equation where the $w_{m}(t)$’s depend on $\psi(t)$ itself, and not a general linear equation. In fact, since we add the preservation of the norm of $\psi(t)$ as a constraint, there exists an equivalent Hermitian rewriting of (29). From the relation $(\mathrm{d}/\mathrm{d}t){\left\langle\psi(t),\psi(t)\right\rangle}=0$, we find that
@@
with the Hermitian operator
@@ VI.2 Geometric principle
For the geometric principle, we have to fix both ${\left\langle\psi,\mathcal{O}_{1}\psi\right\rangle}=\rho_{1}$ and ${\left\langle\psi,\mathcal{O}_{2}\psi\right\rangle}=\rho_{2}=1-\rho_{1}$ to ensure the normalization of $\psi$. Recall that $S^{\psi}$ stays invertible under the sole condition that $0<\rho_{1}(t)<1$ for all times. No other condition is needed. This is an important difference compared with the variational principle, which had the additional constraint (49) on the velocity $\rho^{\prime}_{1}(t)$. With the geometric principle, all time-dependent densities are representable, even those changing very fast. This was already observed in [44, Appendix].
@@
These formulas can also be found in [44, Appendix]. When $\rho_{1}(t)=1/2$ and $\alpha_{2}(0)=\alpha_{1}(0)$ or $\alpha_{2}(0)=\alpha_{1}(0)+\pi$, we find the stationary solutions corresponding to the two eigenstates of $H$. If $\rho_{1}(t)\equiv\rho_{1}$ is time-independent with $\rho_{1}\neq 1/2$ and $\beta(0)$ is not equal to 0 or $\pi$ modulo $2\pi$, then the relative angle $\beta(t)$ will depend on time through (59) and the solution $\psi(t)$ will not be an eigenstate.
@@ VII.2 $v$-representability and the invertibility of the matrices $S^{\Psi}$ and $K^{\Psi}$
Recall that the geometric and variational principles respectively rely on the two matrices $S^{\Psi(t)}$ and $K^{\Psi(t)}(t)$. Namely, we get a well-defined flow as long as these matrices are invertible. In this section we give an interpretation of the invertibility in relation with the Hohenberg–Kohn theorem and the $v$-representability problem. We particularly focus on the initial time $t=0$ and therefore do not emphasize the time-dependence in our notation. We thus simply denote the Hamiltonian by $H_{U}$.
@@
such that $\rho_{\Psi}=\rho$. In principle $E_{0}$ could be any eigenvalue of $H_{U}+\mathcal{V}$ with $\mathcal{V}:=\sum_{m=1}^{M}v_{m}\,\mathcal{N}_{m}$, but we restrict our attention to ground states. Not all densities are $v$-representable in this sense [46, 47, 48, 49, 50, 51, 52, 53]. More generally, we say that a density is ensemble $v$-representable if we can find a mixed state $\Gamma$ supported in the ground eigenspace of $H_{U}+\mathcal{V}$, i.e. $(H_{U}+\mathcal{V}-E_{0})\Gamma=0$, such that $\rho_{\Gamma}=\rho$. If the ground state is non-degenerate (that is, the associated eigenspace has dimension one), then the two are equivalent, of course. It was shown in [50] that any density satisfying $0<\rho_{m}<2$ is ensemble $v$-representable but for pure states the situation is less clear. We refer to [52] for simple examples of densities that are not $v$-representable by a pure state.
@@
This proves that the invertibility of $S^{\Psi}$ is equivalent to that of the derivative $\chi[v]$ of the potential-to-density map. By the implicit function theorem, the latter implies the invertibility of the full potential-to-density map in a neighborhood of $v$ (modulo an additive constant). This gives yet another interpretation of the invertibility of $S^{\Psi}$, for non-degenerate ground states, which is in the spirit of [52, Thm. 13].
@@ VII.6 Time-dependent current-density functional theory
In addition to the density observables $\mathcal{N}_{m}$, we consider also the current-density operators
@@
where $\tau$ is the hopping parameter. As discussed above, we can remove $\mathcal{N}_{M}$ from the list of observables by choosing the gauge of the added potential to have $v_{M}=0$. Similarly, we can remove $\mathcal{J}_{M}$ since the total current is fixed by the time-dependent density. (Of course, $\mathcal{J}_{M}=0$, unless we consider periodic boundary conditions, where we identify $M+1$ with $1$.) Thus, the list of observables to constrain is $\{\mathcal{N}_{m},\mathcal{J}_{m}\}_{m=1}^{M-1}$. For the variational principle, the question is thus to find real numbers $v_{m}(t)$ and $\alpha_{m}(t)$ such that the state $\Psi^{\mathrm{V}}(t)$ solving

### v2

@@ II.2 Geometric structure of the set of constrained states for time-independent constraints
We have to be careful to work with the real structure of our state space $\mathbb{C}^{d}$, and not the complex structure. This is because it is not the same to be differentiable in the real sense or in the complex sense. (Differentiable in this context refers to a function of $\psi$ being differentiable with respect to the components $\psi_{j}$ of $\psi$.) By real structure we mean that we see any vector $\psi$ in $\mathbb{C}^{d}$ as a vector in $\mathbb{R}^{2d}$, each component $\psi_{j}$ being split into its real part $\Re(\psi_{j})$ and imaginary part $\Im(\psi_{j})$. The usual real scalar product in $\mathbb{R}^{2d}$ of two vectors $\psi,\psi^{\prime}\in\mathbb{C}^{d}$ is
@@
The set $\mathcal{C}$ in (4) can be decomposed into a regular part $\mathcal{M}$ and a singular part $\mathcal{S}$ as $\mathcal{C}=\mathcal{M}\cup\mathcal{S}$. The regular part $\mathcal{M}$ is the set where the constraints are non-redundant (or qualified in the language of constrained optimization) in the sense that the $\mathcal{O}_{m}\psi$’s are $\mathbb{R}$-linearly independent. It follows that $\mathcal{M}$ is a smooth (infinitely differentiable) submanifold of $\mathbb{R}^{2d}$ of dimension 2dM2d-M, see for instance [21, Section 7.7] or [22, Theorem 5.12]. The singular part $\mathcal{S}$ is interpreted as a boundary of $\mathcal{M}$, where (some of) the constraints become redundant. We will restrict ourselves to states evolving only within the regular part $\mathcal{M}$.
@@
$\mathcal{T}_{\psi}:=\big\{h\in\mathbb{C}^{d}\ |\ \Re{\left\langle h,\mathcal{O}_{m}\psi\right\rangle}=0,\ \forall m=1,...,M\big\}.$ (5)
@@
We assume that our initial state belongs to the regular part $\psi_{0}\in\mathcal{M}$. That is, we assume that $S^{\psi_{0}}$ is invertible. In addition, we will only work with trajectories $t\mapsto\psi(t)$ defined over an interval of times $t\in[0,T)$ for which $S^{\psi(t)}$ remains invertible. Note that a matrix very similar to the matrix $S^{\psi}$ already appeared in a work on the abstract extension of the Hohenberg–Kohn theorem [18]. For the special case of commuting observables $\mathcal{O}_{m}$’s, the matrix $S^{\psi}$ also appears in a recent work [20] in the context of constrained search in imaginary time for ground-state DFT. The invertibility of $S^{\psi}$ is closely linked to the unique $v$-representability problem [23] in the case of TDDFT, see Section VII.2. We discuss further the invertiblity of $S^{\psi}$ in Appendix C. Generically, one should expect $S^{\psi}$ to be invertible.
@@ II.3 Time-dependent constraints
Note that existence and uniqueness of such a $\nu_{\psi(t)}$ follow from the fact that (11) consists of $M$ linearly independent equations, and that $\mathcal{N}_{\psi(t)}$ is $M$-dimensional. From the definition (6) of $\mathcal{N}_{\psi}$, we can write
@@
Historically, this problem was studied first by Dirac [24] and Frenkel [25, p. 253]. These authors were however considering the manifold of Slater determinants, that has a natural complex structure, so that several of the difficulties we will encounter do not occur in their case. Meyer, Kǔcar and Cederbaum were probably the first to notice in [26] that several definitions of the constrained Schrödinger dynamics that were the same for Dirac and Frenkel can give different results in other situations. The general theory was further developed in [27, 28, 29, 30, 31] but, to the best of our knowledge, it was never applied to the case of constraining expectation values $\left\langle\psi(t),\mathcal{O}_{m}\psi(t)\right\rangle$}. This is the situation of interest for TDDFT. In the next two sections we describe the main two methods used in practice to choose the correction term $F$ so that the solution to the modified Schrödinger equation stays on the considered manifold at all times, and specify the expressions of $F$ in the case of a manifold defined by such constrained expectation values. We will also introduce a new family of constrained Schrödinger equations that interpolate between the previous two methods, which does not seem to have been considered so far.
@@ III.1 Stationarity of the action
In this section, we investigate under which conditions on the observables $\mathcal{O}_{m}$ one can define a unique constrained dynamics solely based on the stationarity of the action. As we will recall below, standard TDDFT fits into this framework by applying the action principle to the wavefunction at fixed density (and not to the density itself as in [3, 33, 34]). Because of this link with standard TDDFT, it is important to understand how the action principle works (or, in fact, why it does not work so easily in this particular case).
@@
The details of this argument can for instance be read in [31, Prop. 3.1]. Stationarity of the action can therefore be re-interpreted as the requirement that the residue of the unconstrained Schrödinger equation $i\partial_{t}\psi(t)-H(t)\psi(t)$ belongs to the normal space of $\psi(t)$ all along the trajectory. This is to make it orthogonal to the small deformations of the path that can only happen in the directions of the tangent space. We emphasize once again the importance of the real structure of the problem in this result due to the appearance of the real scalar product.
@@ III.3 The case of commuting observables
Exactly as before we note that the invertibility of $K^{\psi(t)}(t)$ implies the full-rank condition (7), i.e., the invertibility of $S^{\psi(t)}$, hence the stationarity of the action for the trajectory $\psi(t)$ provided by the theorem. Also, the normalization $\|\psi(t)\|=1$ is automatic and the identity matrix should not be included in the list of the $\mathcal{O}_{m}$’s, otherwise the matrix $K^{\psi(t)}(t)$ cannot be invertible.
@@ IV.1 Geometric Schrödinger equation
A manifold for which the tangent and normal spaces are complex-linear and not just real-linear (i.e., $i\mathcal{T}_{\psi}=\mathcal{T}_{\psi}$ and $i\mathcal{N}_{\psi}=\mathcal{N}_{\psi}$) is called a Kähler manifold. For such a manifold, the variational principle (17) and the geometric principle (28) give the same answer, possibly up to an irrelevant global phase factor. When the manifold is defined by fixing expectation values of some observables $\mathcal{O}_{m}$, the spaces $\mathcal{T}_{\psi}$ and $\mathcal{N}_{\psi}$ will not be complex-linear, however. For instance, if the $\mathcal{O}_{m}$’s are commuting matrices (as they are in TDDFT) then we have $\Im{\left\langle\mathcal{O}_{m}\psi,\mathcal{O}_{n}\psi\right\rangle}=0$ for all $m,n$, as we have seen in the previous section. This implies $\Re{\left\langle i\mathcal{O}_{m}\psi,\mathcal{O}_{n}\psi\right\rangle}=0$, that is $i\mathcal{N}_{\psi}\subset\mathcal{T}_{\psi}$. More abstractly, the expectation value $\left\langle\psi,\mathcal{O}_{m}\psi\right\rangle$} is a function of both $\psi$ and $\psi^{\dagger}$ and hence not a holomorphic function of $\psi$. In particular, the manifold $\mathcal{M}$, which is given by level sets of such functions, cannot be a Kähler manifold.
Since the condition (28) is only based on an orthogonal projection relying on the geometric structure of the constraint manifold $\mathcal{M}$, we call it the geometric principle (GP). In the literature, the latter has been given different names. It is often called the McLachlan principle [36]. The interpretation in terms of a projection on the tangent space was already mentioned by Frenkel in [25, p. 253] and, apparently, in an appendix of the Russian version of the book [24] by Dirac (see the comments in [37]). The McLachlan principle was rediscovered by Nazarov in [38, 39].
@@
Because of the $i$, at first sight it may seem that we are introducing a non-Hermitian perturbation of the Hamiltonian $H(t)$. But this is not the best way to see it. Recall that this is a nonlinear equation where the $w_{m}(t)$’s depend on $\psi(t)$ itself, and not a general linear equation. In fact, since we add the preservation of the norm of $\psi(t)$ as a constraint, we can rewrite (29) in an equivalent form involving a $\psi$-dependent Hermitian Hamiltonian, where the norm-preservation is more apparent. Indeed, from the relation $(\mathrm{d}/\mathrm{d}t){\left\langle\psi(t),\psi(t)\right\rangle}=0$, we find that
@@
with the Hermitian operator (for every given $\psi(t)$)
@@
Recall that the short-time existence of a solution for the variational principle, as given by Theorem 2, relies on the invertibility of the matrix $K^{\psi_{0}}(t=0)$. As we mentioned after Theorem 2, this implies the invertibility of $S^{\psi_{0}}$, so that the solution to the geometric principle also exists by Theorem 3. In this sense, the geometric principle is more robust than the variational principle.
@@ VI.2 Geometric principle
For the geometric principle, we have to fix both ${\left\langle\psi,\mathcal{O}_{1}\psi\right\rangle}=\rho_{1}$ and ${\left\langle\psi,\mathcal{O}_{2}\psi\right\rangle}=\rho_{2}=1-\rho_{1}$ to ensure the normalization of $\psi$. Recall that $S^{\psi}$ stays invertible under the sole condition that $0<\rho_{1}(t)<1$ for all times. No other condition is needed. This is an important difference compared with the variational principle, which had the additional constraint (49) on the velocity $\rho^{\prime}_{1}(t)$. With the geometric principle, all time-dependent densities are representable, even those changing very fast. This was already observed in [49, Appendix].
@@
These formulas can also be found in [49, Appendix]. When $\rho_{1}(t)=1/2$ and $\alpha_{2}(0)=\alpha_{1}(0)$ or $\alpha_{2}(0)=\alpha_{1}(0)+\pi$, we find the stationary solutions corresponding to the two eigenstates of $H$. If $\rho_{1}(t)\equiv\rho_{1}$ is time-independent with $\rho_{1}\neq 1/2$ and $\beta(0)$ is not equal to 0 or $\pi$ modulo $2\pi$, then the relative angle $\beta(t)$ will depend on time through (59) and the solution $\psi(t)$ will not be an eigenstate.
@@ VII.2 $v$-representability and the invertibility of the matrices $S^{\Psi}$ and $K^{\Psi}$
Recall that the geometric and variational principles respectively rely on the two matrices $S^{\Psi(t)}$ and $K^{\Psi(t)}(t)$. Namely, we get a well-defined dynamics as long as these matrices are invertible.
In this section we consider the case when the initial condition $\Psi_{0}$ is such that the matrices $S^{\Psi_{0}}$ and $K^{\Psi_{0}}(t=0)$ are invertible, and give an interpretation of the invertibility in relation with the Hohenberg–Kohn theorem and the $v$-representability problem. The invertibility of the two matrices will then propagate to short times, leading to the local-in-time existence and uniqueness of solutions by Theorems 2 and 3. As we only consider the initial time $t=0$, we drop the time-variable in our notation and denote the Hamiltonian simply by $H_{U}$.
@@
such that $\rho_{\Psi}=\rho$. In principle $E_{0}$ could be any eigenvalue of $H_{U}+\mathcal{V}$ with $\mathcal{V}:=\sum_{m=1}^{M}v_{m}\,\mathcal{N}_{m}$, but we restrict our attention to ground states. Not all densities are $v$-representable in this sense [51, 52, 53, 54, 55, 56, 23, 57, 58]. More generally, we say that a density is ensemble $v$-representable if we can find a mixed state $\Gamma$ supported in the ground eigenspace of $H_{U}+\mathcal{V}$, i.e. $(H_{U}+\mathcal{V}-E_{0})\Gamma=0$, such that $\rho_{\Gamma}=\rho$. If the ground state is non-degenerate (that is, the associated eigenspace has dimension one), then the two are equivalent, of course. It was shown in [55] that any density satisfying $0<\rho_{m}<2$ is ensemble $v$-representable but for pure states the situation is less clear. We refer to [23] for simple examples of densities that are not $v$-representable by a pure state.
@@
This proves that the invertibility of $S^{\Psi}$ is equivalent to that of the derivative $\chi[v]$ of the potential-to-density map. By the implicit function theorem, the latter implies the invertibility of the full potential-to-density map in a neighborhood of $v$ (modulo an additive constant). This gives yet another interpretation of the invertibility of $S^{\Psi}$, for non-degenerate ground states, which is in the spirit of [23, Thm. 13].
@@ VII.6 Time-dependent current-density functional theory
We consider a one-dimensional chain and, in addition to the density observables $\mathcal{N}_{m}$, also the current-density operators
@@
where $\tau$ is the hopping parameter. Note that we label the current operator on the edge $m,m+1$ simply by the index $m$. As discussed above, we can remove $\mathcal{N}_{M}$ from the list of observables by choosing the gauge of the added potential to have $v_{M}=0$. Similarly, we can remove $\mathcal{J}_{M}$ since the total current is fixed by the time-dependent density. (Of course, $\mathcal{J}_{M}=0$, unless we consider periodic boundary conditions, where we identify $M+1$ with $1$.) Thus, the list of observables to constrain is $\{\mathcal{N}_{m},\mathcal{J}_{m}\}_{m=1}^{M-1}$. For the variational principle, the question is thus to find real numbers $v_{m}(t)$ and $\alpha_{m}(t)$ such that the state $\Psi^{\mathrm{V}}(t)$ solving
@@
For a higher-dimensional lattice one can perform a similar analysis, with the difference that the current operators must now carry indices of the edges rather than the sites. The matrix $A^{\Psi_{0}}$ has the same block form, but the matrix $B^{\Psi_{0}}$ is no longer a square matrix and determining when the matrix $A^{\Psi_{0}}$ is invertible is thus slightly more complicated.